OpenStudy (anonymous):

Manuel builds and sells wooden crates. He made a table to show how much his crates cost. Which equation represents the situation? Manuel’s Crates Crates 2 4 6 8 Cost 25 45 65 85 A. y – 2 = −10(x – 25) B. y – 4 = 10(x – 45) C. y – 65 = −10(x – 6) D. y – 85 = 10(x – 8)

OpenStudy (anonymous):

@Umangiasd can you help plz?

OpenStudy (anonymous):

@AcademicGurusInc can you help plz?

OpenStudy (academicgurusinc):

The solution would be let x = number of crates and y= cost, then 10 (2-8) +85 = 25 which confirms that D is the answer.

OpenStudy (academicgurusinc):

We hope this helps. Regards, Academic GurusInc. Twitter (@Academic_Gurus) Facebook (AcademicGurusInc) Youtube (Academic Gurus Inc)

OpenStudy (anonymous):

thank you can you help with more?

OpenStudy (academicgurusinc):

sure

OpenStudy (anonymous):

thanks give me a second plz

OpenStudy (anonymous):

An object moves at a constant rate. The graph shows the object's height at different times. Which equation represents the situation? A. y = −15x + 900 B. y = −15x − 900 C. y = 15x − 900 D. y = 15x + 900

OpenStudy (anonymous):

An object moves at a constant rate. The graph shows the object's height at different times. Which equation represents the situation? A. y = −15x + 900 B. y = −15x − 900 C. y = 15x − 900 D. y = 15x + 900

OpenStudy (anonymous):

hello?

OpenStudy (academicgurusinc):

In this case you need to calculate the slope which is the difference in the y value divided by the difference in the x values. That is -15 then use the linear equation formula y=mx+b to solve for b, which is 900. The equation should be y=-15x +900

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Martin has a gift card to a coffee shop. Each day he buys the same coffee. The table shows the amount remaining on the gift card on different days. How much was on the card on Day 6? $ Martin’s Gift Card Days 2 4 5 7 Amount remaining 24.10 18.20 15.25 9.35

OpenStudy (anonymous):

last one for now

OpenStudy (academicgurusinc):

In this case you, you would apply the same method: Step 1, calculate the slope by dividing the change in the " amount remaining" divided by change in "gift card days" Step 2: Apply linear formula y=mx +b and solve for b (m=slope) Try this to see if you can solve the problem

OpenStudy (academicgurusinc):

Thank you for the medal

OpenStudy (anonymous):

np thanks for th help

OpenStudy (academicgurusinc):

You are welcome. We'd like to feature your question on our Youtube Tutoring Thursday channel so that we can help other students who may find it helpful. Would you be ok with that?

OpenStudy (academicgurusinc):

This is a link to our site: https://www.youtube.com/channel/UCYiI7SmkU4_vhdSzKBWsifg

OpenStudy (academicgurusinc):

Where you able to solve the last question?

OpenStudy (anonymous):

yea sure

OpenStudy (anonymous):

no still confused

OpenStudy (anonymous):

plz help? im confused

OpenStudy (academicgurusinc):

Ok the response video will be posted today and we will notify you as soon as it has been posted. The slope m = ((15.25 -18.20)/(5-4)) = -2.95 Then y= mx + b Now solve b by using any x, y combination, ex. (2, 24.10) this should give you 30 now you know y= -2.95x + 30. Solve for y by replacing x with 6 (for the 6th day)

OpenStudy (academicgurusinc):

In the mean time, your answer should come $12.30.

OpenStudy (academicgurusinc):

I hope that helps.

OpenStudy (anonymous):

thanks for the example

OpenStudy (anonymous):

how did you get 30?

OpenStudy (academicgurusinc):

You are welcome. 24.10 = -2.95 (2) + b. b = 30 .

OpenStudy (anonymous):

@BunnyBree can you help?

OpenStudy (anonymous):

@Umangiasd can you help?

OpenStudy (anonymous):

Martin has a gift card to a coffee shop. Each day he buys the same coffee. The table shows the amount remaining on the gift card on different days. How much was on the card on Day 6? $ Martin’s Gift Card Days 2 4 5 7 Amount remaining 24.10 18.20 15.25 9.35

OpenStudy (anonymous):

plz help

OpenStudy (anonymous):

so confused

OpenStudy (anonymous):

im still confused

OpenStudy (academicgurusinc):

Ok no problem, do you understand how to calculate the slope

OpenStudy (anonymous):

no not really sorry @AcademicGurusInc

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

so when i do the others do i still use -2.95?

OpenStudy (academicgurusinc):

Ok, let's start from the beginning: You have two variables: Gift card days and amount remaining. Lets call gift card days variable "x" and amount remaining variable "y". Next when you are calculating the slope you the formula is m = (y2 -y1) divided by (x2-x1). Where m is the slope and a point is represented by (x, y) so one of the points could be (2, 24.10) Are you following so far?

OpenStudy (academicgurusinc):

yes that is correct

OpenStudy (anonymous):

thank you so much i get it now you r awesome the only thing u don't get is how to get to the 6th place

OpenStudy (anonymous):

how do i get to the 6th place i am confused when i get to the last one

OpenStudy (academicgurusinc):

the reason being that this is a linear relationship. Therefore y= - 2.95x +30 always applies

OpenStudy (anonymous):

i get now what they r saying but i don't get how do i get to day 6 because if skips a day plz help on that

OpenStudy (academicgurusinc):

You can select any day for x and you should be able to calculate how much is remaining (y)

OpenStudy (anonymous):

nvm i get it thanks so much if i need help with more can i mention you?

OpenStudy (anonymous):

so my answer would be 12.30?

OpenStudy (anonymous):

thansk so much

OpenStudy (academicgurusinc):

for the 6th day x=6 so the amount remaining (y) is calculated: y = -2.95 (6) + 30 y = 12.30 That is the amount remaining on the 6th day. Does that make sense?

OpenStudy (anonymous):

THANKS I GOT 100%

OpenStudy (anonymous):

YES I GOT 100% can you help with more?

OpenStudy (academicgurusinc):

Do you understand now?

OpenStudy (academicgurusinc):

Great! You are welcome. Glad we could help. :)

OpenStudy (anonymous):

Solve. 8x = 4x – 32 A. –4 B. 8 C. –8 D. 4

OpenStudy (anonymous):

@AcademicGurusInc can you help with more and yes i understand how thanks

OpenStudy (anonymous):

yes i understand

OpenStudy (anonymous):

i got -8. is that right?

OpenStudy (academicgurusinc):

yes, you are correct. Good job!

OpenStudy (anonymous):

plz help

OpenStudy (anonymous):

i have to get some work done off of k12 before my maw maw gets home

OpenStudy (anonymous):

so can you help?

OpenStudy (academicgurusinc):

In this case start by bringing all "a" variables to the left side of the equation. Subtract 4a from both sides of the equation. Then, to isolate "a", bring the 14 over the to the right side of the equation by subtracting 14 from both sides of the equation. Are you following so far? what do you get?

OpenStudy (academicgurusinc):

You should have -6a = -30 at this stage.

OpenStudy (anonymous):

so i subtract 14-16? and 14-14 and that cancels out? right?

OpenStudy (academicgurusinc):

The next step is to get rid of the -6 in front of the a. You do this by dividing -6 from both sides of the equation. This leaves you with a= 5. Do you understand?

OpenStudy (academicgurusinc):

Yes, that is correct

OpenStudy (anonymous):

so then on my left side i am left with -2a, right?

OpenStudy (anonymous):

16-14=2

OpenStudy (anonymous):

correct?

OpenStudy (anonymous):

then i have to subtract 4a-2?

OpenStudy (anonymous):

how did you get 5?

OpenStudy (academicgurusinc):

It should look like this: 14 - 2a = 4a - 16 14 - 14 -2a = 4a - 16 - 14 -2a = 4a - 30 -2a -4a= 4a -4a - 30 -6a = -30 (-6a)/-6 = -30/-6 a =5

OpenStudy (anonymous):

ok i am going to right that down so give me a second and thank you

OpenStudy (anonymous):

i don't mean to be a bother but how did you get 30 and can you help with more?

OpenStudy (academicgurusinc):

No problem, the " -30" came -16-14, from the second line of by solution above. Do you see that? Unfortunately, I will have to go but I hope this last question makes sense